OpenStudy (anonymous):

Solve the system of equations. Enter your answer as an ordered pair. 3x - 4y = -30 3x + 2y = 6

OpenStudy (anonymous):

by using elimination method when u subtract the second one u will get, -6y = -36 y = -36/-6 y = 6 3x+2(6) = 6 3x+12 = 6 3x = 6-12 x = -6/3 x = -2

OpenStudy (anonymous):

Solve the system of equations. Enter your answer as an ordered pair. 6x + 5y = 61 -5x + 5y = -5

OpenStudy (anonymous):

(6,3)

OpenStudy (anonymous):

you are wrong irishman it is (-2,6)